A circular ring with a radius of 12.00 cm is in a uniform magnetic field of 6.50 x 103 Tesla which points in the positive y direction. calculate the flux where the normal vector to the surface of the ring is:
1
Expert's answer
2020-04-13T10:21:43-0400

The flux


"\Phi=B\cdot A\cdot\cos\alpha=B\cdot\pi R^2\cdot\cos\alpha=0.0065\cdot3.14\cdot0.12^2\cdot\cos0°=0.3mWb" .